OpenStudy (cj1213):

limit of cos x*ln x as x approaches 0+

OpenStudy (anonymous):

what is \[\lim_{x\to 0^+}\cos(x)\]?

OpenStudy (cj1213):

times ln (x)

OpenStudy (anonymous):

i am just asking if you know \[\lim_{x\to 0^+}\cos(x)\]?

OpenStudy (cj1213):

oh um isnt it just 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since cosine is continuous and \(\cos(0)=1\)

OpenStudy (anonymous):

and what is \(\lim_{x\to 0^+}\ln(x)\) ?

OpenStudy (cj1213):

when I did it I got 0 but i wasnt sure if that was right

OpenStudy (anonymous):

oh no it is \(-\infty\)

OpenStudy (anonymous):

so you end up with \(-\infty\)

OpenStudy (cj1213):

oh okay thanks

OpenStudy (anonymous):

yw